# OmniSage: Large Scale, Multi-Entity Heterogeneous Graph Representation Learning

### v2

@@ Abstract.
This paper presents OmniSage, a large-scale representation framework that learns universal representations for a variety of applications at Pinterest. OmniSage integrates graph neural networks with content-based models and user sequence models by employing multiple contrastive learning tasks to effectively process graph data, user sequence data, and content signals. To support the training and inference of OmniSage, we developed an efficient infrastructure capable of supporting Pinterest graphs with billions of nodes. The universal representations generated by OmniSage have significantly enhanced user experiences on Pinterest, leading to an approximate 2.5% increase in sitewide repins (saves) across five applications. This paper highlights the impact of unifying representation learning methods, and we will open source the OmniSage code by the time of publication.
@@
copyright: rightsretainedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Learning latent representationsccs: Computing methodologies Multi-task learning
@@ 1. Introduction
Pinterest is a content discovery platform with over 550 million monthly active users (Pinterest, 2024) engaging with over 1.5 billion Pins every week (Pinterest, 2024). Each Pin consists of an image and text, representing ideas that users can explore and save. Users organize these Pins into boards for ease of access and categorization. Given the vast number of Pins available on Pinterest, it is critical to provide relevant and engaging recommendations for each surface, including the Homefeed, the related Pins feed, and search results. To support Pinterest’s extensive, multidisciplinary applications, it is important to develop a universal, shareable, and rich understanding of entities (Pins, boards, items, queries, and users) (Ying et al., 2018; Baltescu et al., 2022; Agarwal et al., 2024; Pancha et al., 2022).
@@
Representation learning, which involves creating latent embedding vectors for entities, has been crucial in many industrial applications, such as people recommendation (Sankar et al., 2021), news recommendation (Ge et al., 2020), job recommendation (Borisyuk et al., 2024) and shopping (Liu et al., 2022). Representation learning methods have evolved to consume different types of signals. Graph neural networks are specialized in understanding the relationships among entities in graphs (Hamilton et al., 2017b; Kipf and Welling, 2017). Sequence models can capture how users’ interests evolve over time by encoding a sequence of user activities (Pancha et al., 2022). Content-based models convert textual and image content into embeddings (Devlin et al., 2019; He et al., 2022). Despite the individual strengths of these approaches, the development of a method that synergistically integrates the advantages of all these techniques remains relatively unexplored.
@@ 2. Related Work
Graph Neural Networks (GNNs) are a powerful tool to learn representations from the graph data. Several architectures have been developed, including GraphSage (Hamilton et al., 2017b), Graph Convolutional Networks (Kipf and Welling, 2017), Graph Attention Networks (Velickovic et al., 2018) and Graph Transformers (Yun et al., 2019). Our work expands on GraphSage by sampling the neighbors of a given node using an aggregator network. However, there are two differences. Firstly, instead of mean/max/min pooling for aggregation, we utilize a transformer network (Yun et al., 2019). Secondly, we employ random walks (Grover and Leskovec, 2016) to sample neighbors, as it is easy to scale and more effective than uniform sampling (Bojchevski et al., 2020).
Heterogeneous Graphs and GNN Infrastructure: Heterogeneous graphs contain multiple node types and edge types, and they are prevalent in real-world applications. For heterogeneous graphs, many GNN methods (Wang et al., 2019; Yang et al., 2023; Schlichtkrull et al., 2018) use metapaths (Dong et al., 2017), a sequence of node types that define walks, to distinguish different node types.
@@
Non-GNN Representation Learning: While GNNs are a prominent approach, other representation learning methods exist. One such approach leverages content signals, such as text content (Devlin et al., 2019) or visual content (He et al., 2022). Content-based methods are particularly valuable for cold-start cases where limited neighborhood information is available. An alternative approach involves leveraging sequences of items to learn temporal interactions among them. This method is particularly effective for developing user representations because it captures the evolution of user interests over time. For instance, Pinnerformer (Pancha et al., 2022) utilizes transformers to predict future user actions based on sequences of items with which the user has interacted, thereby learning more dynamic user representations. Our work leverages a feature-based contrastive learning task and a sequential prediction task to combine the strengths of the above methods with GNNs.
@@ 3.2. Embedder Architecture
Each node $u$ has a combination of image and text features. We use an in-house pretrained ViT model (Beal et al., 2022) to encode the images. To embed the raw text features, we utilize a simple shared hash-based embedding (Tito Svenstrup et al., 2017) applied to n-gram tokenized text, allowing efficient representation without compromising scalability.
Given a node $u$ of type $t$ and its local neighborhood $N(u,\mathcal{H})$, the features of $u$ and $v\in N(u,\mathcal{H})$ (i.e., all neighbour features) are processed through a MLP and concatenated to form a sequence. This sequence is passed through transformer encoder layers with multi-head self-attention. The output from the last layer corresponding to a learned CLS token is processed through an MLP, followed by $L_{2}$ normalization to yield the final embedding, denoted as $f_{t}(u)$.
@@ 3.3. Objectives
Contrastive learning (Radford et al., 2021; Chen et al., 2020) is a commonly used approach to learning embeddings, as it helps map entities into an embedding space where similar samples (positive pairs) are close together, while dissimilar samples (negative pairs) are far apart. This enables the embeddings to be powerful in downstream applications such as ranking, retrieval, and classification. Our model aims to achieve several objectives with these embeddings. First, we want embeddings of similar nodes (of same or different types) to be close together in the embedding space. Second, the embeddings should effectively encode both the content and the local neighborhood of each node. Lastly, they should capture the sequence of entities in the user’s journey. To achieve these goals, we employ three types of contrastive tasks: entity-entity, entity-feature, and user-entity, as detailed below.
@@ 3.3.1. Entity-Entity Tasks
Graph Edges. Given a node $u$ and its sampled neighbor set $N(u)$, we sample nodes $v\in N(u)$ of type $t$ to construct pairs for the task. Although these pairs can cover all combinations of entity types, we only consider a subset of entity type pairs based on importance and business need. Note that this approach is similar in intention to edge prediction (Hamilton et al., 2017a), but it utilizes a sampled softmax loss instead of the traditional binary cross-entropy loss.
@@ 4.3.1. Pin Representations
Sitewide Repins +0.92% +0.39% +0.30%
Surface Repins +1.22% +0.47% +0.63%
Surface Closeups +0.52% +0.22% +0.40%
Across all surfaces and on different downstream use cases of Pin representations, there are statistically significant improvement in metrics over PinSage. Across these experimental use cases, we observe an increase of over 1.5% in sitewide rePins and over 2% increase in surface-specific repins. Figure 4 shows a visualization of Pins retrieved using OmniSage Pin embeddings.
@@ 4.3.2. Board Representations
We assess the effectiveness of our board representations through online experiments on the Board More Ideas (BMI) surface, which shows Pin recommendations conditioned on a chosen board. Our approach integrates board representations into both the candidate generation (retrieval) and ranking phases. In the retrieval phase, we introduce a specialized candidate generator based on a conditional learned retrieval model (Lin et al., 2024) that leverages Omnisage board embeddings. In the ranking phase, we incorporate board embeddings into the feature mixing layer (Xia et al., 2023). As depicted in Table 8, we observe significant metric improvements over the baselines. Combined across both ranking and retrieval we see improvements of +0.9% in sitewide repins and +5.73 % in BMI repins. Figure 5 shows a visualization of top Pins closest in the embedding space to the OmniSage board embeddings.
@@
Sitewide Repins +0.52% +0.38%
BMI Repins +1.94% +3.79%
BMI Closeups +3.56% +4.65%
@@ A.2.4. Aggregation Technique Evaluation
Though we leverage a simple, out-of-the-box multi-head attention-based aggregation technique across our neighbourhood (and our anchor node’s features), we compare to some existing traditional aggregation techniques to ensure that our technique is optimal. For this purpose, we leverage Graph Attention Networks (GAT), which are reliable and perform strongly across a wide range of use cases (Veličković et al., 2017).

### v3

@@ Abstract.
This paper presents OmniSage, a large-scale representation framework that learns universal representations for a variety of applications at Pinterest. OmniSage integrates graph neural networks with content-based models and user sequence models by employing multiple contrastive learning tasks to effectively process graph data, user sequence data, and content signals. To support the training and inference of OmniSage, we developed an efficient infrastructure capable of supporting Pinterest graphs with billions of nodes. The universal representations generated by OmniSage have significantly enhanced user experiences on Pinterest, leading to an approximate 2.5% increase in sitewide repins (saves) across five applications. This paper highlights the impact of unifying representation learning methods, and we make the model code publicly available  [^1] https://github.com/pinterest/atg-research/tree/main/omnisage.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737253isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Learning latent representationsccs: Computing methodologies Multi-task learning
@@ 1. Introduction
Pinterest is a content discovery platform with over 550 million monthly active users (Pinterest, 2024) engaging with over 1.5 billion Pins every week (Pinterest, 2024). Each Pin consists of an image and text, representing ideas that users can explore and save. Users organize these Pins into boards for ease of access and categorization. Given the vast number of Pins available on Pinterest, it is critical to provide relevant and engaging recommendations for each surface, including the Homefeed, the related Pins feed, and search results. To support Pinterest’s extensive, multidisciplinary applications, it is important to develop a universal, shareable, and rich understanding of entities (Pins, boards, items, queries, and users) (Ying et al., 2018; Baltescu et al., 2022; Agarwal et al., 2024; Pancha et al., 2022; Pal et al., 2020; Yang et al., 2020).
@@
Representation learning, which involves creating latent embedding vectors for entities, has been crucial in many industrial applications, such as people recommendation (Sankar et al., 2021), news recommendation (Ge et al., 2020), job recommendation (Borisyuk et al., 2024) and shopping (Liu et al., 2022; Fan et al., 2019). Representation learning methods have evolved to consume different types of signals. Graph neural networks are specialized in understanding the relationships among entities in graphs (Hamilton et al., 2017; Kipf and Welling, 2017). Sequence models can capture how users’ interests evolve over time by encoding a sequence of user activities (Pancha et al., 2022; Hansen et al., 2020; Zhou et al., 2024). Content-based models convert textual and image content into embeddings (Devlin et al., 2019; He et al., 2022; Radford et al., 2021a; Beal et al., 2022a; Wang et al., 2024). Despite the individual strengths of these approaches, the development of a method that synergistically integrates the advantages of all these techniques remains relatively unexplored.
@@ 2. Related Work
Graph Neural Networks (GNNs) are a powerful tool to learn representations from the graph data. Several architectures have been developed, including GraphSage (Hamilton et al., 2017), Graph Convolutional Networks (Kipf and Welling, 2017), Graph Attention Networks (Veličković et al., 2018) and Graph Transformers (Yun et al., 2019). Our work expands on GraphSage by sampling the neighbors of a given node using an aggregator network. However, there are two differences. Firstly, instead of mean/max/min pooling for aggregation, we utilize a transformer network (Yun et al., 2019). Secondly, we employ random walks (Grover and Leskovec, 2016; Perozzi et al., 2014) to sample neighbors, as it is easy to scale and more effective than uniform sampling (Bojchevski et al., 2020).
Heterogeneous Graphs and GNN Infrastructure: Heterogeneous graphs contain multiple node types and edge types, and they are prevalent in real-world applications. For heterogeneous graphs, many GNN methods (Wang et al., 2019; Yang et al., 2023; Schlichtkrull et al., 2018) use metapaths (Dong et al., 2017; Wang et al., 2023), a sequence of node types that define walks, to distinguish different node types.
@@
Non-GNN Representation Learning: While GNNs are a prominent approach, other representation learning methods exist. One such approach leverages content signals, such as text content (Devlin et al., 2019; Zhuang and Liu, 2019; Liu et al., 2020) or visual content (He et al., 2022; Beal et al., 2022a; Xu et al., 2023). Content-based methods are particularly valuable for cold-start cases where limited neighborhood information is available. An alternative approach involves leveraging sequences of items to learn temporal interactions among them. This method is particularly effective for developing user representations because it captures the evolution of user interests over time. For instance, Pinnerformer (Pancha et al., 2022) utilizes transformers to predict future user actions based on sequences of items with which the user has interacted, thereby learning more dynamic user representations. Our work leverages a feature-based contrastive learning task and a sequential prediction task to combine the strengths of the above methods with GNNs.
@@ 3.2. Embedder Architecture
Each node $u$ has a combination of image and text features. We use an in-house pretrained ViT model (Beal et al., 2022b) to encode the images. To embed the raw text features, we utilize a simple shared hash-based embedding (Svenstrup et al., 2017) applied to n-gram tokenized text, allowing efficient representation without compromising scalability.
Given a node $u$ of type $t$ and its local neighborhood $N(u,\mathcal{H})$, the features of $u$ and $v\in N(u,\mathcal{H})$ (i.e., all neighbour features) we first enforce an ordering on this neighborhood using the proximity scores calculated during the random walk like $\mathrm{Seq}(u)=(v_{1},v_{2},\dots,v_{k})$ such that $s(u,v_{1},\mathcal{H})\geq s(u,v_{2},\mathcal{H})\geq\dots\geq s(u,v_{k},\mathcal{H})$. The neighbor features are then passed through an MLP and concatenated to form a sequence. This sequence is passed through transformer encoder layers with multi-head self-attention. The output from the last layer corresponding to a learned CLS token is processed through an MLP, followed by $L_{2}$ normalization to yield the final embedding, denoted as $f_{t}(u)$.
@@ 3.3. Objectives
Contrastive learning (Radford et al., 2021b; Chen et al., 2020) is a commonly used approach to learning embeddings, as it helps map entities into an embedding space where similar samples (positive pairs) are close together, while dissimilar samples (negative pairs) are far apart. This enables the embeddings to be powerful in downstream applications such as ranking, retrieval, and classification. Our model aims to achieve several objectives with these embeddings. First, we want embeddings of similar nodes (of same or different types) to be close together in the embedding space. Second, the embeddings should effectively encode both the content and the local neighborhood of each node. Lastly, they should capture the sequence of entities in the user’s journey. To achieve these goals, we employ three types of contrastive tasks: entity-entity, entity-feature, and user-entity, as detailed below.
@@ 3.3.1. Entity-Entity Tasks
Graph Edges. Given a node $u$ and its sampled neighbor set $N(u)$, we sample nodes $v\in N(u)$ of type $t$ to construct pairs for the task. Although these pairs can cover all combinations of entity types, we only consider a subset of entity type pairs based on importance and business need. Note that this approach is similar in intention to edge prediction (Hamilton et al., 2017), but it utilizes a sampled softmax loss instead of the traditional binary cross-entropy loss.
@@ 4.3.1. Pin Representations
Sitewide Repins +0.92% +0.39% +0.85%
Surface Repins +1.22% +0.47% +1.35%
Surface Closeups +0.52% +0.22% +1.12%
Across all surfaces and on different downstream use cases of Pin representations, there are statistically significant improvement in metrics over PinSage. Across these experimental use cases, we observe an increase of over 1.7% in sitewide repins and over 3% increase in surface-specific repins. Figure 4 shows a visualization of Pins retrieved using OmniSage Pin embeddings.
@@ 4.3.2. Board Representations
We assess the effectiveness of our board representations through online experiments on the Board More Ideas (BMI) surface, which shows Pin recommendations conditioned on a chosen board. Our approach integrates board representations into both the candidate generation (retrieval) and ranking phases. In the retrieval phase, we introduce a specialized candidate generator based on a conditional learned retrieval model (Lin et al., 2024) that leverages Omnisage board embeddings. In the ranking phase, we incorporate board embeddings into the feature mixing layer (Xia et al., 2023). As depicted in Table 8, we observe significant metric improvements over the baselines. Combined across both ranking and retrieval we see improvements of +1.05% in sitewide repins and +9.7% in BMI repins. Figure 5 shows a visualization of top Pins closest in the embedding space to the OmniSage board embeddings.
@@
Sitewide Repins +0.67% +0.38%
BMI Repins +5.91% +3.79%
BMI Closeups +4.02% +4.65%
@@ A.2.4. Aggregation Technique Evaluation
Though we leverage a simple, out-of-the-box multi-head attention-based aggregation technique across our neighbourhood (and our anchor node’s features), we compare to some existing traditional aggregation techniques to ensure that our technique is optimal. For this purpose, we leverage Graph Attention Networks (GAT), which are reliable and perform strongly across a wide range of use cases (Veličković et al., 2018).